OpenStudy (anonymous):

Okay so just tried to do f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). And i got 5x-3 for my final answer. -Help please

OpenStudy (whpalmer4):

Okay, \[f(x) = -4x+7\]\[g(x) = 10x -6\]You want \(f(g(x))\) The way to think of this is as a recipe. Don't get confused the by fact that we recycle the letter \(x\)! It may be easier to think of this with different letters: \[f(y) = -4y+7\]\[g(x) = 10x-6\] \(g(x)\) is still \(10x-6\) but when we do \(f(g(x))\) we are doing \(f(y) = f(10x-6)\) That means we take \[f(y) = -4y+7\]and replace \(y\) with \((10x-6)\):\[f(y) = f(g(x)) = -4(10x-6) + 7\]\[f(g(x)) = -4*10x -4*(-6) + 7 = -40x +31\] Does that make sense to you?

OpenStudy (anonymous):

im stuck after i times them by each other in paranthases

OpenStudy (whpalmer4):

times what by each other?

OpenStudy (whpalmer4):

You aren't multiplying the functions, you're substituting one function for the argument of the other.

OpenStudy (anonymous):

i get -40x+24-70x-42

OpenStudy (whpalmer4):

Look at my last two lines again: \[f(y) = f(g(x)) = -4(10x-6)+7\] \[f(g(x)) = -4*10x -4*(-6) + 7 = -40x + 24 + 7 = -40x +31\]

OpenStudy (anonymous):

oh god thats confusing

OpenStudy (anonymous):

then i just got 14=14...pellet

OpenStudy (whpalmer4):

Look, \[f(x) = -4x+7\]is just a recipe. It says "take the thing you are given (we'll call it \(x\)), multiply it by -4, and add 7 to the result" So when the thing it is given is the expression \(10x-6\), we multiply \((10x-6) \text{ by } -4\), giving us \(-4(10x-6) = -4*10x -4(-6) = -40x + 24\), and then we add 7 to the result, giving us \(-40x + 24 + 7 = -40x + 31\)

OpenStudy (whpalmer4):

write out your steps for me, one step at a time, so I can point out where you are going astray...

OpenStudy (anonymous):

now i got 40x-17

OpenStudy (anonymous):

geeze

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

i do 4(10x-6)+7 which is 40x-24+7 which then i get 40x-17

OpenStudy (whpalmer4):

No, it's not \(4(10x-6)+7\text{, it's }-4(10x-6)+7\) \[f(x) = -4x+7\]

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

Does that clear it up for you?